# Accelerating Quantum Eigensolver Algorithms With Machine Learning

### v1

@@ Abstract
In this paper, we explore accelerating Hamiltonian ground state energy calculation on NISQ devices. We suggest using search-based methods together with machine learning to accelerate quantum algorithms, exemplified in the Quantum Eigensolver use case. We trained two small models on classically mined data from systems with up to 16 qubits, using XGBoost’s Python regressor. We evaluated our preliminary approach on 20-, 24- and 28-qubit systems by optimising the Eigensolver’s hyperparameters. These models predict hyperparameter values, leading to a 0.13%-0.15% reduction in error when tested on 28-qubit systems. However, due to inconclusive results with 20- and 24-qubit systems, we suggest further examination of the training data based on Hamiltonian characteristics. In future work, we plan to train machine learning models to optimise other aspects or subroutines of quantum algorithm execution beyond its hyperparameters.
@@ 1 Introduction
Modern-day Noisy Intermediate-Scale Quantum (NISQ) devices are the current state-of-the-art in Quantum Computing (QC) characterised as noisy with an intermediate scale in terms of the number of qubits they have [43, 10, 1]. These devices have not yet evolved to support fault-tolerant calculations with a large enough number of qubits to achieve quantum advantage. However, these already enable quantum researchers and engineers to develop, optimise and test various quantum algorithms. In this interim time of intensive hardware development, robust and innovative quantum algorithms [34] provide a way to tackle the limitations of the current quantum processors. One of the most promising applications of QC is in the field of quantum chemistry for molecular simulations, structure design, drug design and more [37, 33, 35]. A prominent example of a class of quantum algorithms that is rapidly evolving in the field of quantum chemistry is the Variational Quantum Eigensolver (VQE) [51], which aims at estimating the ground state energy of Hamiltonians. Different types of VQE algorithms are being developed and tailored for various systems.
@@
Whilst taking part in the Quantum Algorithm Grand Challenge [46] organised by QunaSys [^1] QunaSys is a technology company aiming to achieve QC’s full potential through algorithm development into product engineering [44], we tackle an instance of quantum algorithms for eigensolving, designed to estimate the eigenvalues and eigenvectors of a given Hamiltonian to find its ground state energy. Our approach focuses on applying machine learning algorithms to explore and optimise quantum algorithms to achieve potential advancements in quantum eigensolving and improve the performance and accuracy of quantum algorithms for larger system setups that are currently too complex to tackle. Our first attempt at the problem was to find a methodology to select the best hyperparameters for the ADAPT-QSCI quantum algorithm [32] that minimise the runtime and the final result for the ground state energy through using machine learning techniques trained on smaller systems [13]. This method is not specifically tailored to ADAPT-QSCI and can be applied to any Quantum Eigensolver-based algorithm. Next, we applied the Quantum Complex Exponential Least Squares (QCELS) algorithm [17] to the problem. We investigated the suitability and limitations of the QCELS algorithm when applied to large systems [14]. In the evaluation, we have assessed their performance and potential benefits for solving 16 quantum systems of 20-, 24- and 28-qubits, showing limited minor improvement for 20- and 28-qubit systems.
@@
The rest of the paper is structured as follows. Section 2 lists related work in the field and Section 3 focuses on the background of the quantum computing algorithms that we have investigated, software engineering optimisation methods and machine learning, followed by a deeper dive into the methodologies of quantum eigensolver algorithms in Section 4, and machine learning applications to accelerate quantum algorithms in Section 5. In Section 6, we present the experiment’s methodology and our results. Our conclusions can be found in Section 7, including notes containing references to the source code used in the paper and the relevant data and acknowledgements.
@@
The challenge focused on improving current quantum algorithms for the problem of finding the ground state energy of a one-dimensional orbital rotated Fermi-Hubbard model Hamiltonian using a given 28-qubit system in the most optimal way [46]. The aim of the challenge is to design an algorithm, given specific constraints (i.e. the number of shots of $10^{7}$, timeout of $6\times 10^{5}$ seconds and limited input data), that can output the closest answer to the actual ground state energy. Previous year’s challenge [45] had the same aim but was based on a much smaller 8-qubit system. The winning algorithm was the company’s proposition later published as Quantum-Selected Configuration Interaction (QSCI) [32].
@@ 2 Related Work
Quantum Computing.
Quantum Computing (QC) was conceptualised initially to simulate quantum mechanics using computers “built of quantum mechanical elements which obey [the] quantum mechanical law” [22]. Later, it was found that QC could have several potential applications and offer significant speed-up over classical computing [5, 4, 16, 6, 27, 11]. In 1994, Shor’s proposal of a polynomial-time algorithm for prime factorization and discrete logarithms on a quantum computer raised enormous interest due to its potential threat to modern RSA cryptosystems [49]. Soon after, Grover introduced a fast database search on quantum computers that promised quadratic speed-up over the best classical algorithm [27]. The resulting potential speed-up is often referred to as “quantum supremacy” [3]. Several studies have applied software engineering techniques to optimise quantum computing [24, 25], while testing and debugging approaches have been found to be beneficial in quantum software development [39, 59].
NISQ.
Demonstrating quantum supremacy on real hardware remains a long-standing challenge, especially at a scale where quantum devices would solve real-life calculations. Although quantum supremacy seems difficult to achieve soon, NISQ algorithms are a prominent example that hybrid systems combining small quantum circuits with classical computations could present some computational advantages, i.e., a quantum advantage [43]. Most agree this stage of QC will likely last for the next few years if not decades, and refer to it as the NISQ era [43]. Variational Quantum Algorithms (VQA) are the most common example of an efficient combination of a reduced quantum circuit inside a classical optimisation loop [51].
Machine Learning.
Machine Learning (ML) algorithms are increasingly used to improve and automate software engineering tasks [55, 58, 28, 48, 19], especially after the advent of Large Language Models (LLM), with common applications in software engineering, including optimisation, code generation, bug detection and automated testing [20, 53, 12, 9, 15, 30]. Furthermore, connections between Machine Learning and QC have been broadly explored, both to optimise Machine Learning with QC and to optimise QC with Machine Learning [21, 41, 54]. As we aim to optimise the quantum algorithm hyper-parameters classically, we discuss how we apply machine learning to enhance the software engineering aspects of our approach in Section 5.
@@ 3.1 Variational Quantum Algorithms
    Convergence check and output The optimisation process continues until a convergence criterion is met, indicating that further iterations are unlikely to significantly improve the solution. This convergence check ensures that the algorithm has reached a stable and potentially optimal solution. The final set of optimised parameters $\theta_{opt}$ represents the solution to the quantum problem. This solution can be used for further analysis or as the output of the VQA.
@@
3.2 Use of AI with GI and Search-based Optimisation
In this work, we integrate ML with search-based methods. We first use predictions on smaller systems, utilising the cost function values of a Hamiltonian of smaller systems and randomly generated hyperparameters of the quantum algorithm to train a gradient boosting model to predict the cost function value. We then use this model to explore and optimise hyperparameters for a larger Hamiltonian system, which is costly to solve classically, via a crossover operator.
Search-based Optimisation and Genetic Improvement (GI).
@@ 3.2 Use of AI with GI and Search-based Optimisation
These methods aim to find the near-optimal solution iteratively, ending when a stopping condition is met (e.g. reaching a maximum number of iterations). Starting with a set of candidate solutions, they are evaluated using a fitness function. GI then diversifies solutions over iterations through mutation operations like bit flip (on a single solution) or crossover (on multiple solutions) to potentially generate better-performing offspring. The method can occasionally minimise the population or inject noise to avoid converging to local minima (or maxima, depending on the problem). Using ML to estimate the fitness function is common when direct evaluation or computation is infeasible. We utilise this idea for larger Hamiltonians in Section 5.
@@
In this work: Data consists of Hamiltonians, parameters and the lowest energy levels. We collect data from smaller quantum systems (e.g. 2-qubit, 4-qubit) to build a model for larger systems (20-qubit and above). Data is gathered from online sources or computed via simulations or classical algorithms. The data is then processed and augmented to ensure compatibility with machine learning models, aiming to keep relevant features of the physical system for training. We discuss our choices with respect to data collection, augmentation, and model prediction in Section 5.
@@
The Gradient Boosting technique [23] is a supervised learning method suitable for regression and classification. It efficiently handles large amounts of data, real numbers, and datasets with many features, as is the case here when representing the Hamiltonian in our dataset, which even with aggressive approximation may result in large data instances. The resultant prediction model allows prediction with large datasets with high precision, avoiding flat predictions caused by scaling issues or oversimplified predictions due to insufficient data relative to the number of features.
@@ 4 Quantum Eigensolver Algorithms
In our evaluation, we used the implementation of ADAPT-QSCI provided by the challenge scripts [46] and implemented our own QCELS solver. Each implementation can run on a classical simulator (i.e. classic mode) or a NISQ (i.e. QC mode). In classical mode, we are limited by the size of the system and probably can run up to 20-qubit systems at most. In our experiments, we limited our classical mode runs to 16-qubit systems.
@@ 4.1 ADAPT-QSCI Algorithm
Quantum-selected configuration interaction (QSCI) method is a computational algorithm in quantum chemistry that is used for calculating electronic structure of molecules in an intelligently chosen subspace that makes larger systems feasible to study on modern-day NISQ devices [31]. A full configuration interaction (FCI) requires high computational cost and memory usage that is out of reach for large systems but by using QSCI the computational space can be reduced through selecting only the most important configurations (ways of distributing electrons among molecular orbitals) by a pre-selection algorithm. One such algorithm is ADAPT-QSCI [40].
@@
Adaptive Construction of Input State for Quantum-Selected Configuration Interaction (ADAPT-QSCI) is an iterative algorithm that uses a predetermined pool of single Pauli operators $\text{I\kern-1.49994ptP}=\{P_{1},\dots,P_{T}\}$ that are generators of rotation gates for the input quantum state of QSCI, and selects the best operators from the pool to lower the energy output by QSCI. The operator pools are similar to ADAPT-VQE approaches and could be based on fermionic or qubit excitations. In this method a simple sampling measurement is performed on an input state prepared by a quantum computer - this is the only step that requires quantum computation. The measurement result is used for identifying the most important electron configurations for performing the selected configuration interaction (CI) calculation on classical computers, that is, Hamiltonian diagonalization in the selected $R_{k}$ dimensional subspace $S_{k}=\operatorname{span}\{\ket{r_{1}^{(k)}},\ldots,\ket{r_{R_{k}}^{(k)}}\}$ [40]. QSCI relies on a quantum computer only for generating the electron configurations via sampling, and the subsequent calculations to output the ground-state energy and the pool operator gradients $h_{j}=\braket{c_{k}}{i[H,P_{j}]}{c_{k}}$ are executed on classical computers333|$\ket{c_{k}}$ is a state corresponding to classical vector $c_{k}=\sum_{l=1}^{R_{k}}(c_{k})_{l}\ket{r_{l}^{(k)}}$, and $i[H,P_{j}]$ is calculated through projecting onto the subspace $S_{k}$ and evaluating the expectation classicaly using the classical vector $c_{k}$.[40].. These calculations are made possible on classical machines due to the reduction of the dimensionality of the subspace based on QSCI. The quantum advantage of the ADAPT-QSCI stems from the potential speed up and improved precision of the generating of the electron configurations via sampling.
@@ 4.2 Quantum Complex Exponential Least Squares (QCELS) Algorithm
Figure 1: QCELS circuit [17]
The QCELS algorithm [17] takes a reference state |$\ket{\psi_{0}}$ and evolves it by the time evolution operator $U(t)=e^{-iHt}$ enclosed within a Hadamard test (as depicted by Figure 1). If the reference state is the ground state (|$\ket{\phi_{0}}$) then the resultant expectation values will have a single frequency $Z_{n}=\braket{\phi_{0}}{U(t_{n})}{\phi_{0}}=e^{-iE_{0}t_{n}}$, where $E_{0}$ is the ground state energy. If the reference state, |$\ket{\psi_{0}}$, is not exactly the ground state then the resultant function is $\ Z_{n}\approx\braket{\psi_{0}}{U(t_{n})}{\psi_{0}}=\sum_{i}p_{i}e^{-iE_{i}t_{n}}$ where $p_{i}=|\braket{\phi_{i}}{\psi_{0}}|^{2}$ is the probability of measuring the eigenstate |$\ket{\phi_{i}}$ of the Hamiltonian. Thus, if a reference state with good overlap with the true ground state is known, we can apply the time evolution operator within a Hadamard test $N$ times and fit to the resulting complex exponential. For this challenge, we found that the provided Hamiltonians had an initial Hartree-Fock ground state that retains an overlap with the true ground state - up to $28$ qubits - to extract a good estimate of the energy. The time evolution operator was implemented with a first-order Trotter-Suzuki expansion with Hamiltonian truncated to only the Pauli operators with the largest 200 coefficients to reduce gate count.
@@
which were constrained to be smaller in magnitude than the initial fitted frequency $\theta_{1}^{(1)}$ to ensure that they acted as corrections to the initial fit, for the purpose of the challenge we also added a factor $0.8$ to keep the fitting procedure as safe as possible due to the need for it to be automated. The size of the amplitude $r_{2}^{(2,3)}$ is also constrained to be smaller than $r_{1}^{(1)}$.
@@
Figure 2: QCELS results from sample data of the orbital rotated Hubbard Hamiltonian on 28 qubits (various parameters)

### v2

@@ Abstract
In this paper, we explore accelerating Hamiltonian ground state energy calculation on NISQ devices. We suggest using search-based methods together with machine learning to accelerate quantum algorithms, exemplified in the Quantum Eigensolver use case. We trained two small models on classically mined data from systems with up to 16 qubits, using XGBoost’s Python regressor. We evaluated our preliminary approach on 20-, 24- and 28-qubit systems by optimising the Eigensolver’s hyperparameters. These models predict hyperparameter values, leading to a 0.12% reduction in error when tested on 28-qubit systems. However, due to inconclusive results with 20- and 24-qubit systems, we suggest further examination of the training data based on Hamiltonian characteristics. In future work, we plan to train machine learning models to optimise other aspects or subroutines of quantum algorithm execution beyond its hyperparameters.
@@ 1 Introduction
Modern-day Noisy Intermediate-Scale Quantum (NISQ) devices are the current state-of-the-art in Quantum Computing (QC) characterised as noisy with an intermediate scale in terms of the number of qubits they have [55, 12, 1]. These devices have not yet evolved to support fault-tolerant calculations with a large enough number of qubits to achieve quantum advantage. However, these already enable quantum researchers and engineers to develop, optimise and test various quantum algorithms. In this interim time of intensive hardware development, robust and innovative quantum algorithms [39] provide a way to tackle the limitations of the current quantum processors. One of the most promising applications of QC is in the field of quantum chemistry for molecular simulations, structure design, drug design, and more [45, 38, 43]. A prominent example of a class of quantum algorithms that is rapidly evolving in the field of quantum chemistry is the Variational Quantum Eigensolver (VQE) [67], which aims at estimating the ground state energy of Hamiltonians. Different types of VQE algorithms are being developed and tailored for various systems.
@@
Whilst taking part in the Quantum Algorithm Grand Challenge [58] organised by QunaSys [^1] QunaSys is a technology company aiming to achieve QC’s full potential through algorithm development into product engineering [56], we tackle an instance of quantum algorithms for eigensolving, designed to estimate the eigenvalues and eigenvectors of a given Hamiltonian to find its ground state energy. Our approach focuses on applying machine learning algorithms to explore and optimise quantum algorithms to achieve potential advancements in quantum eigensolving and improve the performance and accuracy of quantum algorithms for larger system setups that are currently too complex to tackle. Our first attempt at the problem was to find a methodology to select the best hyperparameters for the ADAPT-QSCI quantum algorithm [36] that minimises the runtime and the final result for the ground state energy through using machine learning techniques trained on smaller systems [10]. This method is not specifically tailored to ADAPT-QSCI and can be applied to any Quantum Eigensolver-based algorithm. Next, we applied the Quantum Complex Exponential Least Squares (QCELS) algorithm [19] to the problem. We investigated the suitability and limitations of the QCELS algorithm when applied to large systems [40]. In the evaluation, we have assessed their performance and potential benefits for solving 16 quantum systems of 20-, 24- and 28-qubits, showing improvement for some of the 20- and 28-qubit systems.
@@
The rest of the paper is structured as follows. Section 2 lists related work in the fields of quantum computing and machine learning for quantum. Section 3 focuses on the background of the quantum computing algorithms that we have investigated, software engineering optimisation methods and machine learning, followed by a deeper dive into the methodologies of quantum eigensolver algorithms in Section 4, and machine learning applications to accelerate quantum algorithms in Section 5. In Section 6, we present the experiment’s methodology and our results. Our conclusions can be found in Section 7, including notes containing references to the source code used in the paper and the relevant data and acknowledgements.
@@
The challenge focused on improving current quantum algorithms for the problem of finding the ground state energy of a one-dimensional orbital rotated Fermi-Hubbard model Hamiltonian using a given 28-qubit system in the most optimal way [58]. The aim of the challenge is to design an algorithm, given specific constraints (i.e. the number of shots of $10^{7}$, timeout of $6\times 10^{5}$ seconds and limited input data), that can output the closest answer to the actual ground state energy. The 2023 challenge [57] had the same aim but was based on a much smaller 8-qubit system. The winning algorithm was the company’s proposition, later published as Quantum-Selected Configuration Interaction (QSCI) [36].
@@ 2 Related Work
Challenges in Quantum Computing. Quantum Computing (QC) was conceptualised initially to simulate quantum mechanics using computers “built of quantum mechanical elements which obey quantum mechanical law” [26]. Later, it was found that QC could have several potential applications and offer significant speed-up over classical computing [6, 5, 17, 7, 31, 13]. In 1994, Shor’s proposal of a polynomial-time algorithm for prime factorization and discrete logarithms on a quantum computer raised enormous interest due to its potential threat to modern RSA cryptosystems [63]. Soon after, Grover introduced a fast database search on quantum computers [31] that promised quadratic speed-up over the best classical algorithm. The resulting potential speed-up is often referred to as “quantum supremacy” [3]. Several studies apply software engineering techniques to optimise quantum computing [29, 28]. Noticeably, testing [48, 65], debugging [51, 62, 46], verification [76, 41] approaches, and efficient synthesis techniques [35, 70, 52] have been found to be beneficial in quantum software development [48, 65, 51, 76].
Demonstrating quantum supremacy on real hardware remains a long-standing challenge, especially at a scale where quantum devices would solve real-life calculations. Although quantum supremacy seems difficult to achieve soon, NISQ algorithms—Imperfect hardware is often called Noisy Intermediate-Scale Quantum (NISQ) devices—are a prominent example that hybrid systems combining small quantum circuits with classical computations could present some computational advantages, i.e., a quantum advantage [55]. Most agree this stage of QC will likely last for the next few years if not decades, and refer to it as the NISQ era [55]. Variational Quantum Algorithms (VQA) are the most common example of an efficient combination of a reduced quantum circuit inside a classical optimisation loop [67]. Other algorithms use classical optimisation to optimise quantum calculations, such as QCELS [20, 21] that uses a fitting procedure to optimise quantum calculations. Because of their prominent role in modern Quantum Computing research and industrial applications, we chose to focus our study on Variational and QCELS Algorithms.
Machine Learning for Quantum Software Engineering. Machine learning algorithms are increasingly used to improve and automate software engineering tasks [74, 77, 32, 61, 23], especially after the advent of Large Language Models (LLM), with common applications in software engineering, including optimisation, code generation, bug detection and automated testing [24, 72, 14, 11, 16, 34]. In Section 5, we suggest applying these ideas in the context of quantum algorithms, accelerating their performances We apply a search-based process to identify the optimal combination of the hyperparameters (global minima) using our regressor trained on smaller systems to estimate the system energy as the fitness function. We give here an overview of the methods used: search-based methods, the regressor and its Python implementation. [25, 73, 53].
Although existing applications of Machine Learning in Quantum Computing have been investigated [60], it remains a very young and open field of research. While Quantum Machine Learning offers potential speedups in Machine Learning tasks [69], very recent work also demonstrates promising results in applying Machine Learning to Optimise Quantum Computations [15, 42].
@@ 3.1 Variational Quantum Algorithms
    Convergence check and output. The optimisation process continues until a convergence criterion is met, indicating that further iterations are unlikely to significantly improve the solution. This convergence check ensures that the algorithm has reached a stable and potentially optimal solution. The final set of optimised parameters $\theta_{opt}$ represents the solution to the quantum problem. This solution can be used for further analysis or as the output of the VQA.
@@
3.2 Machine Learning and Software Engineering Methods
In this work, we integrate ML with search-based methods. We first use predictions on smaller systems, utilising the cost function values of a Hamiltonian of smaller systems and randomly generated hyperparameters of the quantum algorithm to train a gradient boosting model to predict the cost function value. We then use this model to explore and optimise hyperparameters for a larger Hamiltonian system.
Search-based Software Engineering.
@@ 3.2 Machine Learning and Software Engineering Methods
These methods aim to find the near-optimal solution iteratively, ending when a stopping condition is met (e.g. reaching a maximum number of iterations). Starting with a set of candidate solutions, they are evaluated using a fitness function. In this work, we utilise the genetic algorithms (GA) approach. GA diversifies solutions over iterations through mutation operations like bit flip (on a single solution) or crossover (on multiple solutions) to potentially generate better-performing offspring. The method can occasionally minimise the population or inject noise to avoid converging to local minima (or maxima, depending on the problem). Using ML to estimate the fitness function is common when direct evaluation or computation is infeasible. We utilise this idea for larger Hamiltonians in Section 5.
@@
In this work: Data consists of Hamiltonians, parameters and the lowest energy levels. We collect data from smaller quantum systems (16 qubits and below) to build a model for larger systems (20 qubits and above). Data is gathered from online sources or computed via simulations or classical algorithms. The data is then processed and augmented to ensure compatibility with machine learning models, aiming to keep relevant features of the physical system for training. We discuss our choices with respect to data collection, augmentation, and model prediction in Section 5.
@@
The Gradient Boosting technique [27] is a supervised learning method suitable for regression and classification. Supervised machine learning utilises labelled data to train predictive models [37].
Gradient boosting efficiently handles large amounts of data, real numbers, and datasets with many features, as is the case here when representing the Hamiltonian in our dataset, which, even with aggressive approximation, may result in large data instances. The resultant prediction model allows prediction with large datasets with high precision, avoiding flat predictions caused by scaling issues or oversimplified predictions due to insufficient data relative to the number of features.
@@ 4 Quantum Eigensolver Algorithms
In our evaluation, we used the implementation of ADAPT-QSCI provided by the challenge scripts [58] and implemented our own QCELS solver. Each implementation can run on a classical statevector (exact) simulator (i.e. classic mode) or on a matrix product state (MPS) simulator (i.e. QC mode) [79, 4, 78, 50], provided as part of the challenge, which is suitable for simulating large system sizes by enabling a trade-off between speed and accuracy. In classical mode, we are limited by the size of the system and probably can run up to 20-qubit systems at most. In our experiments, we limited our classical mode runs to 16-qubit systems.
@@ 4.1 ADAPT-QSCI Algorithm
Quantum-selected configuration interaction (QSCI) method is a computational algorithm in quantum chemistry that is used for calculating electronic structure of molecules in an intelligently chosen subspace that makes larger systems feasible to study on modern-day NISQ devices [36]. A full configuration interaction requires high computational cost and memory usage that is out of reach for large systems but by using QSCI the computational space can be reduced through selecting only the most important configurations (ways of distributing electrons among molecular orbitals) by a pre-selection algorithm. One such algorithm is ADAPT-QSCI [49].
@@
Adaptive Construction of Input State for Quantum-Selected Configuration Interaction (ADAPT-QSCI) is an iterative algorithm that uses a predetermined pool of single Pauli operators $\text{I\kern-1.49994ptP}=\{P_{1},\dots,P_{T}\}$ that are generators of rotation gates for the input quantum state of QSCI, and selects the best operators from the pool to lower the energy output by QSCI. The operator pools are similar to ADAPT-VQE approaches and could be based on fermionic or qubit excitations. In this method a simple sampling measurement is performed on an input state prepared by a quantum computer - this is the only step that requires quantum computation. The measurement result is used for identifying the most important electron configurations for performing the selected configuration interaction calculation on classical computers, that is, Hamiltonian diagonalization in the selected $R_{k}$ dimensional subspace $S_{k}=\operatorname{span}\{\ket{r_{1}^{(k)}},\ldots,\ket{r_{R_{k}}^{(k)}}\}$ [49]. QSCI relies on a quantum computer only for generating the electron configurations via sampling, and the subsequent calculations to output the ground-state energy and the pool operator gradients $h_{j}=\braket{c_{k}}{i[H,P_{j}]}{c_{k}}$ are executed on classical computers333|$\ket{c_{k}}$ is a state corresponding to classical vector $c_{k}=\sum_{l=1}^{R_{k}}(c_{k})_{l}\ket{r_{l}^{(k)}}$, and $i[H,P_{j}]$ is calculated through projecting onto the subspace $S_{k}$ and evaluating the expectation classicaly using the classical vector $c_{k}$.[49].. These calculations are made possible on classical machines due to the reduction of the dimensionality of the subspace based on QSCI. The quantum advantage of the ADAPT-QSCI stems from the potential speed up and improved precision of the generating of the electron configurations via sampling.
@@ 4.2 Quantum Complex Exponential Least Squares (QCELS) Algorithm
Figure 1: QCELS circuit [19].
The QCELS algorithm [19] takes a reference state |$\ket{\psi_{0}}$ and evolves it by the time evolution operator $U(t)=e^{-iHt}$ enclosed within a Hadamard test (as depicted by Figure 1). If the reference state is the ground state (|$\ket{\phi_{0}}$) then the resultant expectation values will have a single frequency $Z_{n}=\braket{\phi_{0}}{U(t_{n})}{\phi_{0}}=e^{-iE_{0}t_{n}}$, where $E_{0}$ is the ground state energy. If the reference state, |$\ket{\psi_{0}}$, is not exactly the ground state then the resultant function is $\ Z_{n}\approx\braket{\psi_{0}}{U(t_{n})}{\psi_{0}}=\sum_{i}p_{i}e^{-iE_{i}t_{n}}$ where $p_{i}=|\braket{\phi_{i}}{\psi_{0}}|^{2}$ is the probability of measuring the eigenstate |$\ket{\phi_{i}}$ of the Hamiltonian. Thus, if a reference state with good overlap with the true ground state is known, we can apply the time evolution operator within a Hadamard test $N$ times and fit to the resulting complex exponential. For this challenge, we found that the provided Hamiltonians had an initial Hartree-Fock ground state that retains an overlap with the true ground state - up to $28$ qubits - to extract a good estimate of the energy. The time evolution operator was implemented with a first-order Trotter-Suzuki expansion with Hamiltonian truncated to only the Pauli operators with the largest coefficients to reduce gate count.
@@
which were constrained to be smaller in magnitude than the initial fitted frequency $\theta_{1}^{(1)}$ to ensure that they acted as corrections to the initial fit, for the purpose of the challenge we also added a factor to keep the change in $\theta_{1}$ small ensuring that the fitting procedure is as reliable as possible due to the need for it to be automated. The size of the amplitude $r_{2}^{(2,3)}$ is also constrained to be smaller than $r_{1}^{(1)}$.
@@
Figure 2: QCELS results from sample data of the orbital rotated Hubbard Hamiltonian on 28 qubits (various parameters).